How many molecules are there in 985ml if nitrogen at 0.0°C and 1.00x10⁶mm Hg
1
Expert's answer
2020-11-09T14:07:08-0500

V (N2)= 985 ml = 985x10-6 m3

T = 0.0°C = 273,2 K

p =  1,00x10⁶mm Hg = 133,32x10⁶ Pa

pV = nRT

R = 8.314 J⋅K−1⋅mol−1

n = pV/(RT)

n(N2) = (133,32x10⁶ x 985x10-6) / (8,314*273,2)= 57,815 moles

n = N/NA

N = n x NA

N(N2) = 57,815 x 6.02x1023 = 3.48x1025